# AVID: Any-Length Video Inpainting with Diffusion Model

### 06 Dec 2023 v1

@@ 4.3 Ablation Analysis
Figure 8: Analysis of Temporal MultiDiffusion. While direct sampling of two video clips (w/o) results in inconsistent content within the target area, the temporal multi-diffusion sampling (w/) ensures the synthesis of longer videos with seamless transitions. We highlight part of the target region in each video to better illustrate the effectiveness of our approach.
@@
Temporal MultiDiffusion. We conduct longer-video experiments on $32$-frame $6$-FPS videos, while our base model is trained on video clips consisting of $16$ frames. As in 3.4, we can directly apply our model to the longer video due to the flexibility of our motion modules, but it will results in drastic quality degradation Another naive approach is to apply our model twice on each non-overlapped 16 frames. However, this approach results in a significant change in the identity of the generated content within the target region between the two segments. In  Fig. 8, we demonstrate the advantages of the Temporal MultiDiffusion pipeline we propose. We set attention guidance $\omega=0$ in this experiment to better illustrate the effect of Temporal MultiDiffusion sampling. For illustration, we highlight the $15^{\text{th}}$ frame (the end of the first segment) and the $16^{\text{th}}$ frame (the beginning of the second segment). As observed, there is a sudden change between the segments using the naive approach, e.g., there is a sudden shift in the texture pattern on the generated goose’s neck and the position of the generated raccoon. As comparison, the content transitions smoothly between any two consecutive frames, with the Temporal MultiDiffusion module in place.
@@
Figure 9: Analysis of middle-frame attention guidance. Adjusting $\omega$ under a consistent random seed yields editing results of differing quality.
@@
Effect of middle-frame attention guidance. We delve into the effect of the attention guidance weight, denoted as $\omega$, as illustrated in Fig. 9. Our experimental videos are with $32$ frames and $6$ FPS. Setting $\omega=0$ bypasses our proposed self-attention guidance mechanism. Although Temporal MultiDiffusion sampling ensures a smoother transition of content within the generated area, there persists a noticeable identity shift throughout the video. For instance, with $\omega=0$, the rendered MINI Cooper gradually transitions from red to dark red. Our proposed attention guidance significantly mitigates this issue, ensuring a consistent identity from the first frame to the last. Conversely, setting a too large value for $\omega=1$ may introduce artifacts, as the keys and values in each self-attention layer Eq. 6 are dominated by the middle frame signal.
@@
Discussion. Our model performance is limited by the learned motion module quality. There are scenarios that our model cannot generate content well especially when the editing prompt involves complicated actions, e.g., “transforming the head of a horse from left to right” (in supplementary material). We believe a stronger motion module or equivalently a better text-to-video foundation model could further help improving the inpainting performance. Another promising direction for future exploration is turn the hyper parameter of structure guidance scale into a learnable parameter, controlled by the editing prompt.
@@ 5 Conclusion
We present AVID, a novel approach for addressing text-guided video inpainting. Our model incorporates a motion module for better temporal consistency, and a structure guidance module for better structural fidelity to the original video. At the sampling phase, we introduce a zero-shot inference pipeline, enabling our model to deal with videos with extended lengths. Additionally, we propose a simple middle-frame attention guidance mechanism, greatly improving the identity consistency across video frames. Rigorous experiments show the effectiveness and robustness of AVID.
@@ Appendix A Qualitative Results
To facilitate a more interactive and illustrative experience, these qualitative results are provided in video format. Readers are recommended to check these results in the first section of our accompanying webpage. These visualization provides a more nuanced understanding of the temporal and visual qualities of our video inpainting results, as well as a deeper insight into the effectiveness of AVID  in practical applications.
@@ Appendix B Exploring Additional Inpainting Tasks
This section delves into the adaptability of our AVID  method to a broader spectrum of video inpainting applications, specifically focusing on content removal and environment swapping. Our experiments are designed to illustrate the versatility and effectiveness of AVID  in handling diverse inpainting scenarios.
The experiments in this section are conducted using videos with $N^{\prime}=24$ frames, corresponding to a duration of $4$ seconds. In these experiments, we set $\omega_{s}=0.0$, meaning no structure guidance is applied. T he results are visually represented in  Fig. A1 of the supplementary material.
@@ B.1 Content Removal
The primary goal in content removal is to eliminate a specific object or element from the video while maintaining seamless integration with the surrounding content. As demonstrated in the top block of Fig. A1, our method initiates this process by generating a mask sequence targeting the object to be removed. Subsequently, we input a prompt such as “A wheat field” that describes the desired background, omitting any mention of the target object. This strategy enables our model to effectively remove the object, replacing it with contextually coherent content that blends seamlessly with the surrounding area.
@@ Appendix C Test-time Efficiency
In this section, we engage in comparative experiments with TokenFlow [13], a state-of-the-art video editing method, to demonstrate the effectiveness of our proposed approach. Following the methodology outlined in  Sec. 4.2 in our main paper, we assess the performance of TokenFlow against our method, particularly focusing on tasks such as re-texturing and object swapping. Our evaluation utilizes the same set of videos and automatic metrics detailed in  Sec. 4.2. This comparative study aims to provide an objective and quantifiable measure of each method’s capabilities.
@@
An additional consideration is TokenFlow’s use of DDIM inversion [53] for temporal consistent latent initialization. In contrast, our method employs initialization from a standard Gaussian distribution. This fundamental difference in initialization strategy highlights TokenFlow’s limitations in tasks where no guidance can be obtained from the source video in target region, such as video uncropping.
@@ E.2 Middle-frame Attention Guidance
Figure A3: Ablation analysis of attention guidance. We compare our middle-frame attention guidance approach (MF) with other temporal correlation modeling method variants, including Sparse-Casual Attention (SC_Attn), Middle-frame Sparse-Casual Attention (MSC_Attn) and First-frame attention guidance (FF).
@@
where $Q^{i}=W^{Q}\psi^{i}$, $K^{i}=W^{K}\left[\psi^{1},\psi^{i-1}\right]$, and $V^{i}=W^{V}\left[\psi^{1},\psi^{i-1}\right]$. We adapt Sparse-Casual Attention within each segment of our Temporal MultiDiffusion pipeline.
@@
Results and discussion: Our experiments, visualized in  Fig. A3, demonstrate the varying degrees of success in addressing identity shift issues. The Sparse-Casual Attention (row 2) struggles to prevent identity shifts due to using different key-frame within each segment. Middle-frame Sparse-Casual Attention (row 3) yields better identity preservation, although inconsistencies in the generated patterns can still be observed. The approach using the first frame as guidance (row 4), while maintaining pattern stability, still exhibits significant color variance between the first and last frames.
@@ Appendix F Limitations
As shown in the first case, in an attempt to generate a horse moving its head from left to right, our method fails to generate plausible results. The generated video, instead of showing a smooth head movement, exhibits the head of the horse disappearing and then reappearing on the right side. Concurrently, the body of the horse undergo an unnatural morphing, transitioning from facing left to right with only minor shape changes. Another challenging scenario involves a lion walking forward. The generated video inaccurately shows the left foot of the lion moving through its right foot, an evident deviation from natural movement. For both cases, we recommend viewing the video results for a more comprehensive understanding of these issues.
@@ Appendix G Any-length Text-to-Video Generation
A promising direction for future research lies in the realm of sequential storytelling through video. This involves the idea of performing inference with a series of text prompts, effectively guiding the attention mechanism to evolve in tandem with the narrative. Such an approach could revolutionize the way stories are visually narrated, aligning the generated video content with a progressive textual storyline.

### 20 Mar 2024 v2

@@ 4.3 Ablation Analysis
Figure 8: Analysis of Temporal MultiDiffusion. While direct sampling of two video clips (w/o) results in inconsistent content within the target area, temporal multi-diffusion sampling (w/) ensures the synthesis of longer videos with seamless transitions. We highlight part of the target region in each video to better illustrate the effectiveness of our approach.
@@
Temporal MultiDiffusion. We conduct longer video experiments on $32$-frame $6$-FPS videos, while our base model is trained on video clips consisting of $16$ frames. As in 3.4, we can directly apply our model to the longer video due to the flexibility of our motion modules, but it will result in drastic quality degradation Another naive approach is to apply our model twice on each non-overlapped 16 frames. However, this approach significantly changes the identity of the generated content within the target region between the two segments. In  Fig. 8, we demonstrate the advantages of the Temporal MultiDiffusion pipeline we propose. We set attention guidance $\omega=0$ in this experiment to better illustrate the effect of Temporal MultiDiffusion sampling. For illustration, we highlight the $15^{\text{th}}$ frame (the end of the first segment) and the $16^{\text{th}}$ frame (the beginning of the second segment). As observed, there is a sudden change between the segments using the naive approach, e.g., there is a sudden shift in the texture pattern on the generated goose’s neck and the position of the generated raccoon. In comparison, the content transitions smoothly between any two consecutive frames, with the Temporal MultiDiffusion module.
@@
Figure 9: Analysis of middle-frame attention guidance. Adjusting $\omega$ yields editing results of differing quality.
@@
Effect of middle-frame attention guidance. We delve into the effect of the attention guidance weight, denoted as $\omega$, as illustrated in Fig. 9. Our experimental videos are with $32$ frames and $6$ FPS. Setting $\omega=0$ bypasses our proposed self-attention guidance mechanism. Although Temporal MultiDiffusion sampling ensures a smoother transition of content within the generated area, a noticeable identity shift persists throughout the video. For instance, with $\omega=0$, the rendered MINI Cooper gradually transitions from red to dark red. Our proposed attention guidance significantly mitigates this issue, ensuring a consistent identity from the first frame to the last. Conversely, setting a too large value for $\omega=1$ may introduce artifacts, as the keys and values in each self-attention layer Eq. 6 are dominated by the middle frame signal.
@@
Discussion. Our model performance is limited by the learned motion module quality. There are scenarios in which our model cannot generate content well especially when the editing prompt involves complicated actions, e.g., “transforming the head of a horse from left to right” (in supplementary material). We believe a stronger motion module or a better text-to-video foundation model could further help improve the inpainting performance. Another promising direction for future exploration is to turn the hyper-parameter of the structure guidance scale into a learnable parameter, controlled by the editing prompt.
@@ 5 Conclusion
We present AVID, a novel approach to address text-guided video inpainting. Our model incorporates motion modules for better temporal consistency, and a structure guidance module for better structural fidelity to the original video. At the sampling phase, we introduce a zero-shot inference pipeline, enabling our model to deal with videos with extended lengths. Additionally, we propose a simple middle-frame attention guidance mechanism, greatly improving the identity consistency across video frames. Rigorous experiments show the effectiveness and robustness of AVID.
Acknowledgments: This research has been partially funded by grants to D. Metaxas through NSF: 2310966, 2235405, 2212301, 2003874, and AFOSR-835531.
@@ Appendix A Qualitative Results
To facilitate a more interactive and illustrative experience, these qualitative results are provided in video format. Readers are recommended to check these results in the first section of our accompanying webpage. This visualization provides a more nuanced understanding of the temporal and visual qualities of our video inpainting results, as well as a deeper insight into the effectiveness of AVID  in practical applications.
@@ Appendix B Exploring Additional Inpainting Tasks
This section delves into the adaptability of our AVID  method to a broader spectrum of video inpainting applications, specifically focusing on content removal and environment swapping. Our experiments illustrate the versatility and effectiveness of AVID  in handling diverse inpainting scenarios.
The experiments in this section are conducted using videos with $N^{\prime}=24$ frames, corresponding to a duration of $4$ seconds. We set $\omega_{s}=0.0$ in these experiments, meaning no structure guidance is applied. The results are visually represented in  Fig. A1 of the supplementary material.
@@ B.1 Content Removal
Video inpainting has been narrowly defined as content removal in previous literature [29, 72, 77]. However, with diffusion models, we can enable multiple new inpainting tasks as introduced, which traditional approaches cannot handle. This work focuses mainly on content generation/editing guided by a given prompt and mask. Nevertheless, our model does also support “content removal”.
The primary goal in content removal is to eliminate a specific object or element from the video while maintaining seamless integration with the surrounding content. As demonstrated in the top block of Fig. A1, our method initiates this process by generating a mask sequence targeting the object to be removed. Subsequently, we input a prompt such as “A wheat field” that describes the desired background, omitting any mention of the target object. This strategy enables our model to effectively remove the object, replacing it with contextually coherent content that blends seamlessly with the surrounding area. We further qualitatively evaluate our model on the popular DAVIS [44] dataset to better illustrate the ability of our method, as shown in  Fig. A2.
Figure A2: Content removal on DAVIS [44] dataset. We apply our method for content removal on different videos in the DAVIS [44] dataset. All frames of each video are passed to our model. Frames shown in the figure are evenly distributed in each video. We use prompts “a field”, “a grassland”, and “a park” respectively for these videos.
@@ B.2 Environment Swap
B.3 Multiple Regions Inpainting
Our method is not limited to inpainting one specific region in a video. Independent inpainting can be achieved sequentially for multiple objects. As shown in Fig. A3, we conduct re-texturing on two different regions, i.e. coat, and hair, further demonstrating the effectiveness of our method in real-world applications.
Figure A3: Multiple objects inpainting. We show how our approach can be applied to inpaint multiple objects in a video independently.
@@ Appendix C Test-time Efficiency
In this section, we engage in comparative experiments with TokenFlow [13], a state-of-the-art video editing method, to demonstrate the effectiveness of our proposed approach. We are following the methodology outlined in  Sec. 4.2 in our main paper, we assess the performance of TokenFlow against our method, particularly focusing on tasks such as re-texturing and object swapping. Our evaluation utilizes the same set of videos and automatic metrics detailed in  Sec. 4.2. This comparative study aims to provide an objective and quantifiable measure of each method’s capabilities.
@@
An additional consideration is TokenFlow’s use of DDIM inversion [56] for temporal consistent latent initialization. In contrast, our method employs initialization from a standard Gaussian distribution. This fundamental difference in initialization strategy highlights TokenFlow’s limitations in tasks where no guidance can be obtained from the source video in the target region, such as video uncropping.
@@ E.2 Middle-frame Attention Guidance
Figure A5: Ablation analysis of attention guidance. We compare our middle-frame attention guidance approach (MF) with other temporal correlation modeling method variants, including Sparse-Casual Attention (SC_Attn), Middle-frame Sparse-Casual Attention (MSC_Attn), and First-frame attention guidance (FF).
@@
where $Q^{i}=W^{Q}\psi^{i}$, $K^{i}=W^{K}\left[\psi^{1},\psi^{i-1}\right]$, and $V^{i}=W^{V}\left[\psi^{1},\psi^{i-1}\right]$. A similar technique is also adopted in Pix2Video [7]. We adapt Sparse-Casual Attention within each segment of our Temporal MultiDiffusion pipeline.
@@
Results and discussion: Our experiments, visualized in  Fig. A5, demonstrate the varying degrees of success in addressing identity shift issues. The Sparse-Casual Attention (row 2) struggles to prevent identity shifts due to using different key-frames within each segment. Middle-frame Sparse-Casual Attention (row 3) yields better identity preservation, although inconsistencies in the generated patterns can still be observed. The approach using the first frame as guidance (row 4), while maintaining pattern stability, still exhibits significant color variance between the first and last frames.
@@
E.3 Test-time Masks Accuracy
Due to using random synthetic masks during training, our model is very robust to inaccurate masks. As shown in Fig. A6, our method can successfully inpaint the video following the given text prompt when the mask region is significantly larger than the region size. However, when the mask area can not cover the whole to-be-replaced object, our method will fail to modify the shape of the object due to the preservation of out-of-region details.
Figure A6: Ablation analysis of mask accuracy. We explore the robustness of our method using different mask regions. Here we show $3$ examples using “inaccurately” expanded, accurate, and “inaccurately” eroded masks with the same prompt of “Mini Cooper”.
@@ Appendix F Limitations
As shown in the first case, in an attempt to generate a horse moving its head from left to right, our method fails to generate plausible results. Instead of showing a smooth head movement, the generated video exhibits the head of the horse disappearing and reappearing on the right side. Concurrently, the body of the horse undergoes an unnatural morphing, transitioning from facing left to right with only minor shape changes. Another challenging scenario involves a lion walking forward. The generated video inaccurately shows the left foot of the lion moving through its right foot, an evident deviation from natural movement. For both cases, we recommend viewing the video results for a more comprehensive understanding of these issues.
@@
Besides the limitations discussed above, we admit our model fails at handling discontinuity, especially objects moving out and back to the video. Such an issue could be mitigated with a more deliberate cross-clip attention injection mechanism, which is a critical direction to further improve the robustness.
@@ Appendix G Any-length Text-to-Video Generation
A promising direction for future research lies in the realm of sequential storytelling through video. This involves the idea of performing inference with a series of text prompts, effectively guiding the attention mechanism to evolve in tandem with the narrative. Such an approach could revolutionize how stories are visually narrated, aligning the generated video content with a progressive textual storyline.